Rebel Factory manufactures computer parts that cost​ $20 each to produce. The fixed cost of production is​ $6,000 per week. If the parts sell for​ $30 each, what is the marginal cost of production per part if the factory produces​ 1,000 parts per​ week?
Answer #1

Answer

Marginal Cost is defined as the additional cost required in order to produce one additional uni of output.

Here Fixed cost = 6000, hence whether he produces or not he still has to incur 6000 and hence, Fixed cost has no role in determining Marginal Cost. It is given that In order to produce 1 additional unit of output he has to incur $20. Thus, in order to produce 1 additional unit of output he has to $20. Marginal cost of production per unit = 20 and is constant whatever quantity he produces. Hence, the marginal cost of production per part if the factory produces​ 1,000 parts per​ week is $20.

Note :

Here it is given that F = Fixed Cost = 6000 , V = Variable cost = 20Q where Q quantity of production.

Hence Total Cost(TC) = F + V = 6000 + 20Q

=> Marginal Cost = d(TC)/dQ = d(6000 + 20Q)/dQ = 20 and is constant whatever be the output.

Hence, Using calculus also, the marginal cost of production per part if the factory produces​ 1,000 parts per​ week is $20.